CHIRAL CAVITY RING-DOWN: ABSOLUTE MEASUREMENT OF OPTICAL ROTATION IN GASES AND LIQUIDS WITH SIGNAL REVERSALS

\underline{LYKOURGOS BOUGAS}, G. E. KATSOPRINAKIS, T. P. RAKITZIS, Department of Physics, University of Crete, and Institute of Electronic Structure and Laser, Foundation for Research and Technology-Hellas 71110 Heraklion-Crete, Greece.

We demonstrate the absolute measurement of optical rotation using a four-mirror cavity-ring-down polarimetric setup. A four-mirror bow-tie cavity allows counter-propagating laser beams, for which symmetry is broken with a longitudinal magnetic field acting upon an intracavity magneto-optic window (producing a Faraday rotation θ_{F}). A chiral sample is introduced in one arm of the cavity, producing a chiral rotation φ_{C}. The different symmetry of the two rotations produces a total rotation of (θ_{F} +φ_{C}) for one laser beam, and (θ_{F} +φ_{C}) for the counter-propagating beam. These rotations produce a polarization beating in the cavity ring-down, of frequency ω_{CW} and ω_{CCW} for the clockwise and counter-clockwise beams, respectively. Analysis of the difference (ω_{CW} − ω_{CCW}) yields the chiral rotation angle ±φ_{C}, where the sign of the angle is determined by the sign of the applied magnetic field. Therefore, subtracting the signals using +B⃗ and B⃗ yields 2φ_{C}, allowing the absolute determination of φ_{C}, without needing to remove the sample [1]. We demonstrate these absolute optical rotation measurements for chiral gases (α-pinene) and chiral liquids. This work is a proof-of-principle demonstration of the experimental setup proposed for the measurement of parity nonconserving optical rotation in atomic systems [1].

1 L. Bougas, G. E. Katsoprinakis, W. von Klitzing, J. Sapirstein, and T. P. Rakitzis, Phys. Rev. Lett 108, 210801